# EMMA: Scaling Mobile Manipulation via Egocentric Human Data

### v1

@@ Abstract
Scaling mobile manipulation imitation learning is bottlenecked by expensive mobile robot teleoperation. We present Egocentric Mobile MAnipulation (EMMA), an end-to-end framework training mobile manipulation policies from human mobile manipulation data with static robot data, sidestepping mobile teleoperation. To accomplish this, we co-train human full-body motion data with static robot data. In our experiments across three real-world tasks, EMMA demonstrates comparable performance to baselines trained on teleoperated mobile robot data (Mobile ALOHA), achieving higher or equivalent task performance in full task success. We find that EMMA is able to generalize to new spatial configurations and scenes, and we observe positive performance scaling as we increase the hours of human data, opening new avenues for scalable robotic learning in real-world environments. Details of this project can be found at https://ego-moma.github.io/.
{strip}[Uncaptioned image]
Figure 1: EMMA learns mobile manipulation policies without collecting mobile manipulation teleoperation data. We achieve this through bridging embodiment kinematic gaps and unified co-training of mobile human data and static robot data.
@@ I Introduction
Mobile manipulation has emerged as one of the most challenging problems in robotics due to the dual demands of navigation and manipulation. While recent advances in robot policy learning have demonstrated impressive capabilities in static manipulation, extending these successes to mobile scenarios introduces substantial challenges. The primary obstacle is data scarcity; current approaches tackling mobile manipulation rely on teleoperation frameworks akin to Mobile ALOHA [1] which face scalability limitations. This fundamental bottleneck limits dataset diversity and deployment robustness, particularly in unpredictable real-world settings where robots must navigate novel spatial configurations while maintaining manipulation precision.
Concurrently, recent work in cross-embodiment learning has explored human data as a scalable source of data for training end-to-end robot policies. Critically, human video data is cheap to collect, does not require a physical robot, and can be collected ergonomically via XR wearables [2, 3]. Human video data has been leveraged to train better visual representations [4], extract high-level object affordance [5, 6, 7, 8], or even as direct action supervision via co-training [2, 3]. However, these works have predominantly focused on table-top manipulation.
To this end, we introduce Egocentric Mobile Manipulation (EMMA), an end-to-end system to train mobile manipulation policies solely from robot static-manipulation and human mobile-manipulation data. While the robot data is captured via teleoperation, the human data is captured simply by wearing Project Aria glasses [9], thus avoiding costly mobile manipulation teleoperation. EMMA represents a step toward a new data paradigm in robot learning, where we imbue a robot with new skills by combining a set of robot teleoperation data with a more diverse pool of human demonstrations. In this work, our robot teleoperation data contains no mobile manipulation demonstrations, and we show that this skill can be effectively transferred from human data.
@@
We evaluate EMMA across three real-world mobile manipulation tasks: table service, handover wine, and grocery shopping. Our experiments reveal three key findings. First, EMMA achieves superior performance compared to baselines trained on teleoperated mobile robot data, showing that human demonstrations can replace costly mobile teleoperation. Second, we observe favorable scaling properties—each additional portion of human data yields greater performance than an equivalent portion of teleoperated robot data. Third, EMMA exhibits robust generalization performance, successfully transferring navigation and manipulation skills to novel environments seen only in human demonstrations, unlike teleoperation-based approaches that fail to adapt. These results open new avenues for scalable robotic learning in real-world environments.
@@ II Related Work
Behavior Cloning (BC). Behavior Cloning (BC) has emerged as an effective approach for robot learning, where policies are trained with direct supervised learning from expert demonstrations. Recent advances have shown remarkable results [10, 11, 12, 13, 14, 15], including the promise of building general-purpose policies by learning from large-scale datasets [11, 15, 16]. In particular, research around large multi-embodiment datasets such as Open-X [15] represents a significant milestone, demonstrating how models trained with diverse robot embodiments can acquire generalizable skills across tasks without task-specific engineering. However, most of these approaches focus on static manipulation tasks, with mobile manipulation remaining a significant challenge.
@@
Learning for Mobile Manipulation. Building on successes in static manipulation, recent works have explored learning-based approaches for mobile manipulation that include skill primitives [17, 18, 19], reinforcement learning with decomposed action spaces  [20, 21, 22], and whole-body control objectives [23, 24, 25]. Unlike these approaches, end-to-end imitation learning enables mapping raw pixel information to whole-body actions, showing promising results through large-scale training [26, 27, 28, 29, 11]. A critical challenge is to collect high-quality mobile manipulation demonstrations. Pioneering works [30, 31] have developed new teleoperation systems to facilitate data collection, including tethered leader-follower system [31], VR headsets [32, 33], motion-capture suits [34, 35, 36], smartphone-based control [37], kinesthetic teaching [38], and full-body teleoperation for humanoid robots [31] . However, despite these advances, collecting high-quality mobile manipulation data for diverse scenarios at scale remains a challenge. We propose to leverage egocentric human data (data collected with first-person perspective) captured by wearable devices as an alternative for scaling up imitation learning for mobile manipulation.
@@
Robot Learning from Human Data. Recent work has focused on two complementary themes—leveraging human videos to bootstrap robot learning and finetuning with reinforcement learning for robust policies. In manipulation, co-training on paired egocentric human and robot demonstrations has been shown to boost skill performance [2], while zero- or few-shot transfer from human videos can be enabled by image inpainting and motion-track priors [39, 40]. Building on these, hierarchical planners extract latent action sequences from humans and distill them via retargeters into whole-body controllers  [5, 31, 41]. In navigation, fusing imitation with RL—by turning keypoint matches into rewards or bootstrapping from behavior cloning—yields more reliable policies  [42, 43], and inverse-dynamics models can pseudo-label passive egocentric video to distill intent-conditioned affordance subroutines [44]. We aim to train mobile manipulation policies from human data in a unified learning framework.
@@ III Hardware and Data Preliminaries
Egocentric Data Collection. In this work, we leverage a wearable smart glass Meta Project Aria [9] as our main data collection platform. Echoing prior work [2], we believe Aria glasses are ideal for capturing egocentric human data due to their ergonomic design and machine perception capability provided by the Machine Perception Service (MPS). Specifically, we leverage Aria glasses to capture both exteroception (wide-FOV egocentric RGB images) and proprioception (hand tracking and global localization) data in human mobile manipulation behaviors.
@@
Figure 2: Left: Architecture of joint human-robot policy learning framework. Our model processes heterogeneous human and robot data through stems and decodes them through various action heads. The navigation head is deployed on the robot during evaluation, demonstrating transfer without robot supervision. Right: Our custom low-cost bimanual mobile manipulator.
@@
Low-cost Bimanual Mobile Manipulator. To effectively utilize egocentric human data for mobile manipulation, the robot hardware platform must resemble human sizes and kinematic workspaces. Drawing inspiration from the “Eve” robot platform introduced in EgoMimic [2], we develop a low-cost custom mobile manipulator that comprises of two 6-DoF ViperX 300s mounted in an identical inverted configuration on a height-adjustable rig. The rig is mounted on an AgileX TRACER differential drive AGV platform, which is capable of moving up to 2m/s. The full system stands a maximum height of 1.75m. Similar to Eve, we propose to leverage Aria glasses as the main egocentric perception sensor for the robot and mount it in a way that emulates the hand-eye configuration of a human adult. This mitigates the human-robot camera device gap and reduces the sensor-manipulator kinematic gap. Each arm is equipped with an Intel Realsense D405 on its wrist to facilitate precise near-range manipulation.
@@
Human and Robot Data Streams. We collect distinct data streams from both human demonstrations ($\mathcal{D}_{H}$) and robot execution ($\mathcal{D}_{R}$). A key shared data stream is the egocentric RGB image $I_{\text{ego}}$, generated by the Aria glasses worn by the human or mounted on the robot. The robot provides RGB streams from its two wrist cameras, $I_{\text{wrist}}$. Proprioceptive data streams capture the state of each embodiment: For the human, we leverage the Aria Machine Perception Service (MPS) [45] to estimate bimanual 3D hand poses $H_{p}\in SE(3)\times SE(3)$ and the 3D head pose $H_{d}\in SE(3)$ (in a SLAM-based world frame). For the robot, proprioception includes the state of its bimanual arms via joint positions $R_{q}\in\mathbb{R}^{2\times 7}$ (including gripper state) and the corresponding end-effector poses $R_{p}\in SE(3)\times SE(3)$.
For navigation, we process the human head pose $H_{d}$ to extract 2D base pose representations. Specifically, we project the 3D head pose onto the ground plane to obtain $h_{\text{base}}^{t}=(x^{t},y^{t},\theta^{t})\in SE(2)$, where $(x^{t},y^{t})$ represents the 2D position and $\theta^{t}$ represents the yaw angle at time $t$. We maintain a displacement-based waypoint history $\mathcal{W}_{t}=\{h_{\text{base}}^{t-k_{i}}\}_{i=1}^{K_{h}}$, where $K_{h}$ is the maximum number of historical waypoints. New waypoints are added when the displacement $\|h_{\text{base}}^{t}-h_{\text{base}}^{t-k_{i}}\|_{2}\geq d_{\text{thresh}}$ (e.g., every 0.5m). This displacement-based sampling ensures consistent spatial resolution regardless of movement speed. For policy learning, these waypoints are transformed into the current egocentric frame as $\tilde{\mathcal{W}}_{t}=\{T_{ego}^{-1}\cdot h_{\text{base}}^{t-k_{i}}\}_{i=1}^{K_{h}}$, where $T_{ego}$ is the transformation from world to current egocentric coordinates. These egocentric waypoints provide speed-invariant spatial context for navigation action prediction.
@@ IV EMMA: System and Algorithm
Figure 3: Given a 3D pose trajectory from a human demonstration, we optimize Eq. 1 to compute a feasible 2D trajectory for a differential drive robot. This resulting trajectory can be directly executed by the robot or used as input for policy learning.
@@ IV-A Data Retargeting and Alignment
A fundamental challenge in leveraging human data ($\mathcal{D}_{H}$) for robot learning lies in the significant embodiment gap, impacting both navigation and manipulation. Humans navigate omnidirectionally with decoupled head-gaze, whereas our robot uses a differential-drive base with kinematically constrained movements. Similarly, human hand motions ($H_{p}$) captured egocentrically differ kinematically and distributionally from robot end-effector motions ($R_{p}$). To enable effective knowledge transfer and co-training in an end-to-end imitation learning setting (Sec. IV-B), we introduce distinct retargeting and alignment strategies for navigation and manipulation.
Bridging Navigation Kinematic Gap. The primary focus of our retargeting efforts is translating human navigation trajectories into commands suitable for our differential-drive robot (Fig. 3). During human demonstrations, we extract 2D base poses $h_{\text{base}}^{t}=(x^{t},y^{t},\theta^{t})$ by projecting the 3D head pose $H_{d}^{t}$ onto the ground plane. These poses form navigation waypoints that capture the human’s intended path.
@@
and constraints $v_{\min}\leq v_{k}\leq v_{\max}$, $\omega_{\min}\leq\omega_{k}\leq\omega_{\max}$.
Here, $p_{k}(\mathbf{z})=(x_{k},y_{k})$ represents the robot position at step $k$ resulting from applying the velocity sequence $\mathbf{z}$, $p_{k}^{d}=(x_{k}^{d},y_{k}^{d})$ is the desired position from the human waypoint, and $\text{wrap}(\cdot)$ ensures angular differences are in $\pi,\pi]$. The weights $\lambda_{\text{pos}}$, $\lambda_{\text{yaw}}$, and $\lambda_{\text{smooth}}$ balance position tracking, heading alignment, and velocity smoothness respectively. This constrained optimization yields smooth, kinematically feasible trajectories that approximate human navigation patterns while respecting differential-drive constraints.
Aligning Manipulation Action Data. For manipulation, we address mismatches between human hand data ($H_{p}$) and robot end-effector data ($R_{p}$). Inspired by prior work like EgoMimic [2], we first unify coordinate frames by transforming all upper-body action chunks (both human and robot) into the reference frame of the camera at the time of observation, using SLAM estimates for human data and hand-eye calibration for robot data. This makes predictions relative to the current view. Second, acknowledging persistent distributional gaps due to biomechanics and sensors, we apply Z-score normalization independently to the transformed pose and action data within each source ($D_{H}$ and $D_{R}$), using their respective dataset statistics.

### v2

@@ Abstract
Scaling mobile manipulation imitation learning is bottlenecked by expensive mobile robot teleoperation. We present Egocentric Mobile MAnipulation (EMMA), an end-to-end framework training mobile manipulation policies from human mobile manipulation data with static robot data, sidestepping mobile teleoperation. To accomplish this, we co-train human full-body motion data with static robot data. In our experiments across three real-world tasks, EMMA demonstrates comparable performance to baselines trained on teleoperated mobile robot data (Mobile ALOHA), achieving higher or equivalent task performance in full task success. We find that EMMA is able to generalize to new spatial configurations and scenes, and we observe positive performance scaling as we increase the hours of human data, opening new avenues for scalable robotic learning in real-world environments. Details of this project can be found at: https://ego-moma.github.io/
@@ I Introduction
Mobile manipulation has emerged as one of the most challenging problems in robotics due to the dual demands of navigation and manipulation. While recent advances in robot policy learning have demonstrated impressive capabilities in static manipulation, extending these successes to mobile scenarios introduces substantial challenges. The primary obstacle is data scarcity; current approaches tackling mobile manipulation rely on teleoperation frameworks akin to Mobile ALOHA [fu2024mobile] which face scalability limitations. This fundamental bottleneck limits dataset diversity and deployment robustness, particularly in unpredictable real-world settings where robots must navigate novel spatial configurations while maintaining manipulation precision.
Concurrently, recent work in cross-embodiment learning has explored human data as a scalable source of data for training end-to-end robot policies. Critically, human video data is cheap to collect, does not require a physical robot, and can be collected ergonomically via XR wearables [egomimic, qiu2025humanoidpolicyhuman]. Human video data has been leveraged to train better visual representations [nair2022r3muniversalvisualrepresentation], extract high-level object affordance [wang2023mimicplay, xu2024flowcrossdomainmanipulationinterface, wen2024anypointtrajectorymodelingpolicy, bharadhwaj2024track2actpredictingpointtracks], or even as direct action supervision via co-training [egomimic, qiu2025humanoidpolicyhuman]. However, these works have predominantly focused on table-top manipulation.
To this end, we introduce Egocentric Mobile Manipulation (EMMA), an end-to-end system to train mobile manipulation policies solely from robot static-manipulation and human mobile-manipulation data. While the robot data is captured via teleoperation, the human data is captured simply by wearing Project Aria glasses [engel2023projectarianewtool], thus avoiding costly mobile manipulation teleoperation. EMMA represents a step toward a new data paradigm in robot learning, where we imbue a robot with new skills by combining a set of robot teleoperation data with a more diverse pool of human demonstrations. In this work, our robot teleoperation data contains no mobile manipulation demonstrations, and we show that this skill can be effectively transferred from human data.
@@
We evaluate EMMA across four real-world mobile manipulation tasks: table service, handover wine, grocery shopping and push chair. Our experiments reveal three key findings. First, EMMA achieves superior performance compared to baselines trained on teleoperated mobile robot data, showing that human demonstrations can replace costly mobile teleoperation. Second, we observe favorable scaling properties—each additional portion of human data yields greater performance than an equivalent portion of teleoperated robot data. Third, EMMA exhibits robust generalization performance, successfully transferring navigation and manipulation skills to novel environments seen only in human demonstrations. These results open new avenues for scalable robotic learning in real-world environments.
@@ II Related Work
Behavior Cloning (BC). Behavior Cloning (BC) has emerged as an effective approach for robot learning, where policies are trained with direct supervised learning from expert demonstrations. Recent advances have shown remarkable results [rt22023arxiv, kim24openvla, black2410pi0, chi2024diffusionpolicy, zhao2023learning, o2024open], including the promise of building general-purpose policies by learning from large-scale datasets [kim24openvla, o2024open, khazatsky2024droid]. In particular, research around large multi-embodiment datasets such as Open-X [o2024open] represents a significant milestone, demonstrating how models trained with diverse robot embodiments can acquire generalizable skills across tasks without task-specific engineering. However, most of these approaches focus on static manipulation tasks, with mobile manipulation remaining a significant challenge.
@@
Learning for Mobile Manipulation. Building on successes in static manipulation, recent works have explored learning-based approaches for mobile manipulation that include skill primitives [sun2022fully, wu2023m, wu2023tidybot, kuang2025skillblender], reinforcement learning with decomposed action spaces  [gu2022multi, yokoyama2023asc, xia2021relmogen, xiong2024adaptive], and whole-body control objectives [fu2023deep, yang2024harmonic, hu2023causal, liu2024opt2skill, ben2025homie, wang2024hypermotion]. Unlike these approaches, end-to-end imitation learning enables mapping raw pixel information to whole-body actions, showing promising results through large-scale training [ahn2022can, brohan2022rt, shafiullah2023bringing, pi0.5, kim24openvla, moo2023arxiv]. A critical challenge is to collect high-quality mobile manipulation demonstrations. Pioneering works [jiang2025behavior, fu2024humanplus] have developed new teleoperation systems to facilitate data collection, including tethered leader-follower system [fu2024humanplus], VR headsets [chen2024arcapcollectinghighqualityhuman, lu2025mobiletelevisionpredictivemotionpriors], motion-capture suits [cisneros2022team, dafarra2024icub3, wholebodygeomretarget], smartphone-based control [pmlr-v164-wong22a], kinesthetic teaching [MOMO-FORCE], and full-body teleoperation for humanoid robots [fu2024humanplus]. However, despite these advances, collecting high-quality mobile manipulation data for diverse scenarios at scale remains a challenge. We propose to leverage egocentric human data (data collected with first-person perspective) captured by wearable devices as an alternative for scaling up imitation learning for mobile manipulation.
@@
Robot Learning from Human Data. Recent work has focused on two complementary themes—leveraging human videos to bootstrap robot learning and finetuning with reinforcement learning for robust policies. In manipulation, co-training on paired egocentric human and robot demonstrations has been shown to boost skill performance [egomimic], while zero- or few-shot transfer from human videos can be enabled by image inpainting and motion-track priors [lepert2025phantomtrainingrobotsrobots, ren2025motion]. Building on these, hierarchical planners extract latent action sequences from humans and distill them via retargeters into whole-body controllers [wang2023mimicplay, fu2024humanplus, ImitationNet]. In navigation, fusing imitation with RL—by turning keypoint matches into rewards or bootstrapping from behavior cloning—yields more reliable policies  [karnan2022voila, ramrakhya2023pirlnav], and inverse-dynamics models can pseudo-label passive egocentric video to distill intent-conditioned affordance subroutines [navsubroutine_egocentricvideos]. We aim to train mobile manipulation policies from human data in a unified learning framework.
@@ III Hardware and Data Preliminaries
Egocentric Data Collection. In this work, we leverage a wearable smart glass Meta Project Aria [engel2023projectarianewtool] as our main data collection platform. Echoing prior work [egomimic], we believe Aria glasses are ideal for capturing egocentric human data due to their ergonomic design and machine perception capability provided by the Machine Perception Service (MPS). Specifically, we leverage Aria glasses to capture both exteroception (wide-FOV egocentric RGB images) and proprioception (hand tracking and global localization) data in human mobile manipulation behaviors.
@@
Figure 2: Left: Architecture of joint human-robot policy learning framework, built on top of [wang2024hpt]. Our model processes heterogeneous human and robot data through stems and decodes them through various action heads. The navigation head is deployed on the robot during evaluation, demonstrating transfer without robot supervision. Right: Our custom low-cost bimanual mobile manipulator.
@@
Low-cost Bimanual Mobile Manipulator. To effectively utilize egocentric human data for mobile manipulation, the robot hardware platform must resemble human sizes and kinematic workspaces. Drawing inspiration from the “Eve” robot platform introduced in EgoMimic [egomimic], we develop a low-cost custom mobile manipulator that comprises of two 6-DoF ViperX 300s mounted in an identical inverted configuration on a height-adjustable rig. The rig is mounted on an AgileX TRACER differential drive AGV platform, which is capable of moving up to 2m/s. The full system stands a maximum height of 1.75m. Similar to Eve, we propose to leverage Aria glasses as the main egocentric perception sensor for the robot and mount it in a way that emulates the hand-eye configuration of a human adult. This mitigates the human-robot camera device gap and reduces the sensor-manipulator kinematic gap. Each arm is equipped with an Intel Realsense D405 on its wrist to facilitate precise near-range manipulation.
@@
Human and Robot Data Streams. We collect distinct data streams from both human demonstrations ($\mathcal{D}_{H}$) and robot execution ($\mathcal{D}_{R}$). A key shared data stream is the egocentric RGB image $I_{\text{ego}}$, generated by the Aria glasses worn by the human or mounted on the robot. The robot provides RGB streams from its two wrist cameras, $I_{\text{wrist}}$. Proprioceptive data streams capture the state of each embodiment: For the human, we leverage the Aria Machine Perception Service (MPS) [aria_mps] to estimate bimanual 3D hand poses $H_{p}\in SE(3)\times SE(3)$ and the 3D head pose $H_{d}\in SE(3)$ (in a SLAM-based world frame). For the robot, proprioception includes the state of its bimanual arms via joint positions $R_{q}\in\mathbb{R}^{2\times 7}$ (including gripper state) and the corresponding end-effector poses $R_{p}\in SE(3)\times SE(3)$. These raw streams serve as the inputs to our system; their processing and transformation into policy inputs and action labels is a core algorithmic contribution detailed in Sec. IV-A.
@@ IV-A Data Retargeting and Alignment
A fundamental challenge in leveraging human data ($\mathcal{D}_{H}$) for robot learning is the embodiment gap, defined as the mismatch in physical morphology and movement kinematics between human and robot embodiments, which affects navigation and manipulation. Humans navigate omnidirectionally with decoupled head-gaze, whereas our robot uses a differential-drive base with kinematically constrained movements. Similarly, human hand motions ($H_{p}$) captured egocentrically differ kinematically and distributionally from robot end-effector motions ($R_{p}$). To enable effective knowledge transfer and co-training in an end-to-end imitation learning setting (Sec. IV-B), we introduce two distinct data processing strategies: optimization-based retargeting for navigation and coordinate-space alignment for manipulation.
Bridging Navigation Kinematic Gap. We process the raw human head pose $H_{d}$ (Sec. IV) to generate both the ground-truth action labels for the policy and the contextual state input.
1. Navigation Action Labels: The primary focus of our retargeting efforts is translating human navigation trajectories into ground-truth action labels suitable for our differential-drive robot (Fig. 3). During human demonstrations, we extract 2D base poses $h_{\text{base}}^{t}=(x^{t},y^{t},\theta^{t})$ by projecting the 3D head pose $H_{d}^{t}$ onto the ground plane. These poses form navigation waypoints that capture the human’s intended path.
Figure 3: Given the ground-plane projection of the human head trajectory, we optimize Eq. 1 to produce a smooth and executable path for our differential-drive robot that can be run directly and used as input for policy learning.
@@
$\operatorname{wrap}(\alpha)=\big((\alpha+\pi)\bmod 2\pi\big)-\pi\in[-\pi,\pi).$ (5)
and constraints $1.6\leq v_{k}\leq 1.6$ m/s and $1.5\leq\omega_{k}\leq 1.5$ rad/s. Here, $p_{k}(\mathbf{z})=(x_{k},y_{k})$ represents the robot position at step $k$ resulting from applying the velocity sequence $\mathbf{z}$, $p_{k}^{d}=(x_{k}^{d},y_{k}^{d})$ is the desired position from the human waypoint, and $\text{wrap}(\cdot)$ ensures angular differences are in $\pi,\pi]$. The weights $\lambda_{\text{pos}}=25$, $\lambda_{\text{yaw}}=2$, and $\lambda_{\text{smooth}}=1$ balance position tracking, heading alignment, and velocity smoothness respectively. This constrained optimization yields smooth, kinematically feasible trajectories that approximate human navigation patterns under differential-drive constraints.
2. Navigation Context: To provide the policy with speed-invariant context for predicting navigation actions, we process the human head pose $H_{d}$ into a history of waypoints by projecting $H_{d}$ onto the ground plane to obtain $h_{\text{base}}^{t}=(x^{t},y^{t},\theta^{t})\in SE(2)$. We then maintain a displacement-based waypoint history $\mathcal{W}_{t}=\{h_{\text{base}}^{t-k_{i}}\}_{i=1}^{K_{h}}$, where $K_{h}$ is the maximum number of historical waypoints. New waypoints are added when the displacement $\|h_{\text{base}}^{t}-h_{\text{base}}^{t-k_{i}}\|_{2}\geq d_{\text{thresh}}$ (e.g., every 0.5m). This displacement-based sampling ensures consistent spatial resolution. For policy learning, these waypoints are transformed into the current egocentric frame as $\tilde{\mathcal{W}}_{t}=\{T_{ego}^{-1}\cdot h_{\text{base}}^{t-k_{i}}\}_{i=1}^{K_{h}}$, where $T_{ego}$ is the transformation from world to current egocentric coordinates. This egocentric waypoint history serves as a contextual input to the policy.
Aligning Manipulation Action Data. For manipulation, we address mismatches between human hand data ($H_{p}$) and robot end-effector data ($R_{p}$). Inspired by prior work like EgoMimic [egomimic], we first unify coordinate frames by transforming all upper-body action chunks (both human and robot) into the reference frame of the camera at the time of observation, using SLAM estimates for human data and hand-eye calibration for robot data. This makes predictions relative to the current view. Second, acknowledging persistent distributional gaps due to biomechanics and sensors, we apply Z-score normalization independently to the transformed pose and action data within each source ($\mathcal{D}_{H}$ and $\mathcal{D}_{R}$), using their respective dataset statistics.